# SplatOverflow: Asynchronous Hardware Troubleshooting

### v2

@@ 5.3.1. Selection and View Control
When a part from the CAD model is selected, it is rendered to the user with a yellow wireframe. This allows the user to view details from the splat while examining the geometry of an individual CAD component. This allows the CAD model to act as a highlight on the splat, indicating the region of the hardware that corresponds to a distinct part.
SplatOverflow allows users to select parts and navigate subassemblies of the CAD model to make sense of occluded parts of the hardware not visible in the splat. Given that the CAD model is aligned to the splat, users can ”peer through” sub-assemblies and view internal components that aren’t visible. This is demonstrated in Figure 12(a).
Remote users can use SplatOverflow scenes to independently explore novel viewpoints within the local user’s environment. This is enabled by the splat component of a SplatOverflow scene. From approximately one minute of video footage, SplatOverflow constructs a 3D Gaussian Spalt (Kerbl et al., 2023) of the local user’s scene. After the splat is aligned onto the CAD model, it occupies the same coordinate space and can be navigated with the same controls as a CAD environment. As a result, remote maintainers do not need to actively guide users to capture parts on their hardware. Instead, users can capture rough videos, and the Gaussian Splat can recover geometry and texture details.
@@ 5.3.2. Gestures
Users communicate in SplatOverflow via gestures. Gestures can be used to request information from the user and instruct actions to be applied to the hardware. Gestures are organized chronologically in an interactive timeline. The timeline elements trigger visualizations of the gesture in the SplatOverflow scene and house user discussions about the gesture.
Gestures can be used to request more information about the scene from the user. This can be done by requesting video annotations or updated splats using the request gesture. When a request gesture is authored, a timeline element appears with a QR code. The QR code lets the user populate the request via SplatOverflow’s mobile app. When the request is populated, the timeline element is linked to the new media submitted by the user. This enables SplatOverflow issues to capture the changing state of a local user’s hardware throughout troubleshooting. An example of the request gesture is shown in Figures 6 and 7.
Gestures can also communicate instructions and actions a user should take on their hardware. Currently, SplatOverflow supports a few gestures meant to convey different actions. These action gestures are all anchored to parts in the CAD model, which are highlighted whenever a gesture is selected in the timeline alongside each gesture’s animation. Each gesture creates a timeline element that contains an additional text description explaining the gesture.
SpaltOverflow’s pointing gestures indicate where on a part in CAD a user should inspect. These gestures are rendered with large green arrows pointing to or away from the CAD component. This lets them communicate different meanings depending on the context, such as tightening or loosening bolts, as shown in Figure 12(b). SplatOverflow’s look-here gesture prompts the user to inspect a specific CAD part from a prescribed perspective. When selected, the CAD part is highlighted, and the viewport shifts to re-orient the local user. SplatOverflow’s move gesture indicates to local users how parts should be manipulated. When selected, the referenced CAD part animates to tween between its original position and the final position defined by the remote maintainer. Finally, SplatOverflow supports a probe gesture for use in workflows involving PCBs. The probe gesture indicates where to place multimeter probes and the reading a user should expect to see.
@@ 5.3.3. Timeline Elements and Discussions
Timeline elements refer to action or request gestures authored by different users. The timeline element contains additional text to clarify what an action or request entails. Users can clarify or follow up on a gesture by replying to the corresponding element in the timeline. This facilitates the asynchronous back-and-forth to help make sense of a given gesture.
@@ 5.4. Indexing Data to a CAD Model
SplatOverflow uses the CAD model to index and query data. This includes technical documentation referencing parts in the CAD model and past SplatOverflow issues. In the case of existing technical documentation, SplatOverflow renders web pages that reference CAD components as illustrated in Figure 12(c). Similarly, SplatOverflow can query the history of past issues to retrieve troubleshooting instructions that reference specific CAD components. This ability to feed structured data from troubleshooting issues into a database that can be queried allows SplatOverflow to become more valuable to hardware communities over time.
@@ 6. Implementation
As shown in Figure 13, SplatOverflow is divided into four sections: Scene Capture, Scene Generation, the SplatOverflow front-end, and the back-end Coordination Server.
The mobile capture interface runs in a browser and is written in JavaScript using the MediaStream Recording API. We typically use it from a smartphone browser. On the server, SplatOverflow uses a C++ 3D Gaussian Splatting implementation by MrNeRF (Kerbl et al., 2023) and the COLMAP package for Structure-from-Motion (SfM) to generate the scene. To align a CAD model to the scan, SplatOverflow uses the method from ArUco SFM Scale Adjustment (Meyer, 2023) and an implementation of Arun’s method (Arun et al., 1987).
The back-end coordination server is a Flask application and SQL database that manages the data associated with a SplatOverflow scene. The SfM and Gaussian Splatting pipelines are run on a PC i7-14700K with 32GB RAM DDR5 and Nvidia 4080S GPU with 24GB of VRAM. On average, a 60-second video takes 122 seconds to generate a scene trained on 360p footage and 250 seconds to generate a scene trained on 1080p footage.
7. Limitations
7.1. Acquiring CAD Models
SplatOverflow requires access to the hardware’s CAD model to populate the scene. Moreover, for SplatOverflow scenes to be useful, the CAD models must be complete. CAD models like this can be challenging to come by outside of open-source hardware projects, which can limit the adoption of SplatOverflow. Moreover, even when CAD files are available, the structure of sub-assemblies, parts, and components is not standardized. As a result, preparing CAD files for use in SplatOverflow can require coercing the structure of the CAD assembly into a compatible shape.
7.2. Structure References in Technical Documentation
Another limitation of SplatOverflow lies in its ability to reference a corpus of existing technical documentation. With the Lumen v3, each page of technical documentation contained links referencing the related parts. This greatly simplified populating the documentation interface with instructions to search through. However, not all hardware will share technical documentation in the same structured fields as Opulo does with Lumen v3. As a result, existing technical documentation may have to undergo post-processing to add the relevant structured data indicating which CAD components are referenced.
7.3. Errors in Tag Placement
@@ 8. Demonstrative Examples
8.1. Verifying the Correct Assembly of the Open Book
SplatOverflow is compatible with non-mechanical CAD designs. The Open Book (Castillo, [n. d.]) e-reader is an open source e-reader that consumers assemble on their own to demystify consumer electronics. Successful assembly is a function of correct and precise soldering. In this example, we show how the maintainer of the Open Book can help users verify correct assembly and track down potential errors that inhibit successful booting using a workflow authored and shared in SplatOverflow. Specifically, the maintainer defines a set of probe points to be checked with a multimeter that can indicate the exact probe points on the PCB the user must test, along with the expected outcome of each test.
The local user starts by creating a SplatOverflow scene to follow the assembly verification guide created by a maintainer. The Open Book has a 20mm tag that is used to align the CAD model of the PCB generated from KiCAD (Team, [n. d.]). In this example, we applied the tag to the PCB, but in practice, such a tag could be printed directly onto the PCB’s silkscreen. Once the SplatOverflow scene is generated, the local user loads the instructions authored by the maintainer into their scene and sees the instructions projected onto their hardware. Figure 14(a) shows the instructions rendered onto their PCB.
The local user references verification instructions loaded into their SplatOverflow scene. As shown in Figure 14(a), instructions are overlaid onto the pins the user must probe. The user follows probing instructions to check for continuity, voltage and resistance at points across the PCB, as shown in Figure 14(b). At each step, the timeline event indicates the expected result of the probing for the user to compare against. The guide authored in a SplatOverflow scene in this application can be contextualized to each local user’s hardware. This contrasts most online guides, which are rendered using virtual objects or images and videos of a different hardware instance.
@@ 8.2. Sharing and Updating Routine Maintenance Workflows on the Prusa MK3S 3D Printer
Asynchronous sharing enables feedback loops, leading to improved workflows. The Prusa MK3S 3D Printer is a popular personal fabrication machine targetting hobbyists. Like all 3D printers and fabrication machines, the printer requires routine maintenance to ensure proper function. A typical maintenance task on the machine is tightening the belts that drive two axes of motion. Poorly tightened belts can lead to failed prints or unintended artifacts in otherwise successful prints. Learning to carry out such tasks is essential to owning and operating personal fabrication machines (Subbaraman and Peek, 2023a). Here, we illustrate how SplatOverflow can be used to share maintenance workflows and how end-user feedback can improve the quality of such workflows.
A local user sees ghosting artifacts on their print and is pointed to a SplatOverflow issue explaining how to adjust the machine’s belt tension. Figure 14(b) shows the workflow rendered in the local user’s workspace. The user can follow the instructions but is unsure how much to tighten the belts. They leave a comment asking the maintainer how they know when to stop tightening the belt. The maintainer realizes that the instruction is underspecified and shares an online resource that uses a microphone to validate belt tension. Here, we show a simple example of how the open and asynchronous nature of SplatOverflow issues allows communities of users to iterate on workflows together.
@@ 8.3. Guided Disassembly of the Billy Bookshelf
SplatOverflow enhances static assemblies like furniture by adding a sticker. The Billy Bookshelf is a mass-produced piece of flat-pack furniture. Ikea supports DIY assembly through easy-to-understand assembly instructions. We demonstrate how SplatOverflow can guide users through step-by-step disassembly.
The Billy is roughly 80 inches tall, making it difficult to scan from all angles; particularly from the top down. The local user places an arUco marker on the back-right corner of the bookshelf and scans the Billy. As the bookshelf is upright during the scan, there is a noticeable degradation of quality above the bookshelf. However, since SplatOverflow aligns the CAD model in the scene, users can infer geometry while still receiving context cues from the scan.
The disassembly workflow is reasonably straightforward. The maintainer (in this case, IKEA) provided a SplatOverflow scene with the correct sequence of opening the cam locks that hold this bookshelf together. Users follow these instructions and sequentially disassemble the shelf. For more complex disassembly, it is possible to have the user update the splats to receive new instructions. This would resemble the scene update we described in the walkthrough. This example shows relevance beyond hobbyist hardware and how adding a sticker to ordinary furniture can enable new collaborative workflows for everyday objects.
@@ 9.1. Study Design
Our study is divided into two parts. First, we examine whether non-expert users can capture their hardware to create a SplatOverflow scene. We examine whether participants were able to generate a scene successfully. Second, we present them with two common issues experienced by users of this hardware. These issues were sourced based on the prevalence of online guides related to fixing them. We evaluate whether users could successfully follow the instructions in their SplatOverflow scene, and ask them to assess the usability of our system.
@@ 9.2. Part 1: Generating a SplatOverflow Scene
Each participant was provided a smartphone running SplatOverflow’s capture interface as a web app. The participants were asked to create a splat of the 3D printer infront of them by recording a video of the hardware from a variety of angles. All participants were able to capture a splat, and 83.3% of participants could successfully generate a SplatOverflow scene. For 33.3% participants, we were able to generate a SplatOverflow scene but did not have enough information in the user-captured video to solve for hardware constraints, i.e. align the degrees of freedom in the hardware. Figure 15 shows how many participants could generate a complete SplatOverflow scene with examples of high and low-quality scenes. We then showed participants the generated scene and explained how they could navigate our interface. Following this the participants filled out a survey evaluating the usability of the scanning interface.
We used their generated SplatOverflow scene for each participant in part 2 of our study. If the user’s input failed to generate a scene (p1, p4, p9, p12), we assigned them to a placeholder scene that captured the same hardware.
@@ 9.3. Part 2: Following Instructions in SplatOverflow
In part 2, we presented the participant with two hardware issues in random order and asked them to follow instructions rendered in SplatOverflow to remedy the issues. The issues were: correcting the belt tension in the printer’s x-axis (issue 1) and clearing a jam in the extruder (issue 2). These issues were selected as they are commonly occuring issues with this printer and have a variety of online tutorials dedicated to them (Repair, 2019; 3D, 2023). We randomize the order of these issues, and asked the participants to follow the instructions to the best of their ability. 91.7% of participants were able to follow the instructions to fix the belt-tension issue and 91.7% of participants could fix the extruder jam issue. Following each issue we asked participants to answer some questions about their experience using SplatOverflow to troubleshoot the issue, as shown in Figure 16. Overall, users found that instructions rendered in SplatOverflow were easy to follow and well-aligned with the scanned hardware.
@@ 9.4. System Usability
After both parts of the study we asked participants to evaluate the overall usability of the system in a survey. We use the System Usability Scale (Brooke, [n. d.]) and our system scored 86.25 / 100.
In addition participants reflected on how this system could be extended to help them in their lives. Participants shared that the benefits of SplatOverflow when facing issues with unfamiliar hardware (p1: It is very helpful for me to use and repair, especially when im not familiar with the machine itself) and they also compared our system to existing methods of troubleshooting (p3: I was working with robots (Turtlebot 4 setup). There documentation was too wordy and it is hard to locate the solution from that. So, we have to post GitHub issues and wait for their response. It was sometime hard to explain them my specific problem. Of course, if there was a system that I could have scanned my problem and send them, it would have saved a lot of time).
@@ 10.1.1. Hardware Support for End-Users
SplatOverflow offers a new way to collaborate on hardware. For workflows such as assembly, routine maintenance, operating instructions, and troubleshooting, SplatOverflow provides a novel way for users to receive guidance from collaborators or retrieve associated documentation. By connecting physical instances of hardware to their CAD model and associated documentation, end-users can inspect the ’source’ of the hardware they are working with. This direct connection between the physical part and documentation is a more direct way to navigate documentation that can exist in disparate locations.
Moreover, users can use SplatOverflow to seek help, analogous to asking for help in a forum. In software, platforms such as GitHub (Dabbish et al., 2012) and StackOverflow (Anderson et al., 2012) have been essential to help users learn about new software, troubleshoot issues and scale maintenance efforts needed to support projects. By supporting asynchronous communication about hardware, SplatOverflow presents a similar workflow for hardware.

### v3

@@ 5.3. Interacting with a SplatOverflow Scene
Figure 10. This figure has three boxes in it: A, B, and C. The left subfigure A shows the outlines of the internal components of a subassembly being rendered in the splat. The middle subfigure B shows an arrow point towards a bolt in a subassembly, indicating the bolt be tightened. The right subfigure C shows the user querying documentation on how to tension the belts on the right side of the machine.
@@ 5.3.1. Selection and View Control
SplatOverflow allows users to select parts and navigate subassemblies of the CAD model to compare the as-designed to the as-built and make sense of occluded parts of the hardware not visible in the splat. When a part from the CAD model is selected, it is rendered to the user with a yellow wireframe. This allows the user to compare details from the splat to the geometry of an individual CAD component. Moreover, the visibility of CAD parts in a subassembly can be controlled, allowing users to ”peer through” sub-assemblies and visualize internal components that are not visible, as shown in Figure 10(a).
Remote users can use SplatOverflow scenes to independently explore novel viewpoints within the local user’s environment. This is enabled by the splat component of a SplatOverflow scene. Once the splat is aligned onto the CAD model, it occupies the same coordinate space and can be navigated with the same controls as a CAD environment.
@@ 5.3.2. Gestures
Users communicate in SplatOverflow via gestures. A gesture is an instruction or request that references an element in the SplatOverflow scene. Gestures are organized chronologically in an interactive timeline. The timeline elements trigger visualizations of the gesture in the SplatOverflow scene and house user discussions about the gesture.
Maintainers can use gestures to request more information from the user. This can be done by requesting video annotations or updated splats using the request gesture. When a request gesture is authored, a timeline element appears with a QR code. The QR code lets the user populate the request via SplatOverflow’s mobile interface and appends the requested information to the scene. Figure 5 shows an example of the request gesture.
Gestures can also communicate instructions or actions a user should take on their hardware. These gestures are all anchored to parts in the CAD model, which are highlighted whenever a gesture is selected in the timeline alongside each gesture’s animation. SpaltOverflow’s pointing gestures indicate where on a part in CAD a user should inspect or attend to. When selected, the referenced CAD part is highlighted, and the viewport shifts to re-orient the local user. SplatOverflow’s move gesture indicates to local users how parts should be manipulated. When selected, the referenced CAD part animates to tween between its original position and the final position defined by the remote maintainer. Finally, SplatOverflow includes a set of action gestures for common troubleshooting operations, such as tightening or loosening specific bolts and probing particular pads on a PCB. When selected, the gestures overlay arrows onto the scene to visualize the operation.
@@ 5.3.3. Timeline Elements and Discussions
Timeline elements refer to gestures authored by different users. The timeline element contains additional text to clarify what a gesture entails. Users can ask for clarification or follow up on a gesture by replying to the corresponding element in the timeline. This facilitates the asynchronous back-and-forth communication to help make sense of a given gesture.
@@ 5.4. Indexing Data to a CAD Model
SplatOverflow uses the CAD model to index and query data. This includes technical documentation referencing parts in the CAD model and past SplatOverflow issues. For existing technical documentation, SplatOverflow renders web pages that reference CAD components, as illustrated in Figure 10(c). Similarly, SplatOverflow can query the history of past issues to retrieve troubleshooting instructions that reference specific CAD components. This ability to feed structured data from troubleshooting issues into a database that can be queried allows SplatOverflow to become increasingly useful to hardware communities over time.
@@ 6. Implementation
Our implementation of SplatOverflow is divided into four sections: Scene Capture, Scene Generation, the SplatOverflow front-end, and the back-end Coordination Server.
The mobile capture interface runs in a browser and is written in JavaScript using the MediaStream Recording API. It is designed to be accessible from modern smartphone browsers. On the server, SplatOverflow uses a C++ 3D Gaussian Splatting implementation by MrNeRF (Kerbl et al., 2023) and the COLMAP package for Structure-from-Motion (SfM) to generate the scene. To align a CAD model to the scan, SplatOverflow uses the method from ArUco SFM Scale Adjustment (Meyer, 2023) and an implementation of Arun’s method (Arun et al., 1987).
The back-end coordination server is a Flask application and SQL database that manages the data associated with a SplatOverflow scene. The SfM (Schönberger and Frahm, 2016) and Gaussian Splatting (Kerbl et al., 2023) pipelines run on a PC equipped with an Intel i7-14700K, 32GB RAM DDR5 and Nvidia 4080S GPU with 24GB of VRAM. On average, a 60-second video takes 122 seconds to generate a scene trained on 360p footage and 250 seconds to generate a scene trained on 1080p footage.
@@ 7. Demonstrative Examples
Figure 11. This figure has three boxes in it: A, B, and C. The left upper subfigure A shows a PCB with a sequence of instructions showing where a user should probe to validate their assembly and an image of a user performing the probing. The left lower subfigure B shows a workflow to adjust the tension of the belts on a 3D printer. It also shows a sequence of replies that add information to the original instruction. The right figure C shows an ikea bookshelf alongside instructions on how to disassemble it.
@@
7.1. Verifying the Correct Assembly of the Open Book E-Reader
SplatOverflow is compatible with non-mechanical CAD designs. The Open Book (Castillo, [n. d.]) e-reader is an open-source e-reader that consumers assemble on their own, with the goal of demystifying consumer electronics. Successful assembly is a function of correct and precise soldering. In this example, we show how the maintainer of the Open Book can help users verify correct assembly and track down potential errors that inhibit successful booting using a workflow authored and shared in SplatOverflow. Specifically, the maintainer defines a set of probe points to be checked with a multimeter that can indicate the exact probe points on the PCB the user must test, along with the expected outcome of each test.
The local user starts by creating a SplatOverflow scene to follow the assembly verification guide created by a maintainer. The Open Book has a 20mm tag that aligns the CAD model of the PCB generated from KiCAD (Team, [n. d.]). In this example, we applied the tag to the PCB, but in practice, such a tag could be printed as part of the PCB’s silkscreen. Once the SplatOverflow scene is generated, the local user loads the instructions authored by the maintainer into their scene and sees them projected onto their hardware. Figure 11(a) shows the instructions rendered onto their PCB.
The local user references verification instructions loaded into their SplatOverflow scene. As shown in Figure 11(a), instructions are overlaid onto the pins the user must probe. The user follows probing instructions to check for continuity, voltage, and resistance at points across the PCB. The guide authored in a SplatOverflow scene can be recontextualized to each local user’s hardware. This contrasts most online guides, which are rendered using virtual objects or images and videos of a different hardware instance.
@@ 7.2. Sharing and Updating Routine Maintenance Workflows on the Prusa MK3S 3D Printer
Asynchronous sharing enables feedback loops, leading to improved workflows. The Prusa MK3S 3D Printer is a popular personal fabrication machine targetting hobbyists. Like all 3D printers and fabrication machines, the printer requires routine maintenance to ensure proper function. A typical maintenance task is tightening the belts that drive two axes of motion. Poorly tightened belts can lead to failed prints or unintended artifacts in otherwise successful prints. Learning to carry out such tasks is essential to owning and operating personal fabrication machines (Subbaraman and Peek, 2023a). We illustrate how SplatOverflow can be used to share maintenance workflows and how end-user feedback can improve the quality of such workflows.
A local user sees ghosting artifacts on their print and is pointed to a SplatOverflow issue explaining how to adjust the machine’s belt tension. Figure 11(b) shows the workflow rendered in the local user’s workspace. The user follows the instructions but is unsure how much to tighten the belts. They leave a comment asking the maintainer how they know when to stop tightening the belt. The maintainer realizes that the instruction is underspecified and shares an online resource that uses a microphone to validate belt tension. Here, we show a simple example of how the open and asynchronous nature of SplatOverflow issues allows communities of users to iterate on workflows together.
@@ 7.3. Guided Disassembly of the Billy Bookshelf
SplatOverflow enhances static assemblies like furniture by adding a sticker. The Billy Bookshelf is a mass-produced piece of flat-pack furniture. IKEA provides DIY assembly through easy-to-understand assembly instructions. We demonstrate how SplatOverflow can guide users through step-by-step disassembly.
The Billy is roughly 80 inches tall, making it difficult to scan from all angles; particularly from the top down. The local user places an ArUco marker on the back-right corner of the bookshelf and scans the Billy. As the bookshelf is upright during the scan, there is a noticeable degradation of quality above the bookshelf. However, because SplatOverflow aligns the CAD model in the scene, users can infer geometry while still receiving context cues from the scan.
The disassembly workflow is reasonably straightforward. The maintainer (in this case, IKEA) provides a SplatOverflow scene with the correct sequence of opening the cam locks that hold this bookshelf together. Users follow these instructions and sequentially disassemble the shelf. For more complex disassembly, users can update the splats to receive new instructions, similar to the scene update described in the walkthrough. This example demonstrates how SplatOverflow can enable new collaborative workflows for mass-produced hardware.
@@ 8.1. Study Design
Our study is divided into two parts. First, we examine whether non-expert users can capture their hardware to create a SplatOverflow scene. Second, we present them with two common issues experienced by users of this hardware. These issues were sourced based on the prevalence of online guides addressing them. We examine whether users can successfully follow instructions through a SplatOverflow scene and assess the usability of our system.
@@ 8.2. Part 1: Generating a SplatOverflow Scene
Figure 12. Has two subfigures in it A and B. The left subfigure A shows six renderings of SplatOverflow scenes. Each corresponds to a user-generated splat in a user study. The splats vary in quality with worst on the left and best on the right. The right subfigure B shows a table of study results.
Participants were provided a smartphone running SplatOverflow’s capture interface as a web app. They were instructed to create a splat of the 3D printer in front of them by recording a video of the hardware from a variety of angles. All participants were able to generate a splat from their recording, and 83.3% of participants could successfully generate a SplatOverflow scene. For 33.3% of participants, the system generated a SplatOverflow scene, but the user-captured video lacked enough information to resolve hardware constraints—specifically, aligning the degrees of freedom in the hardware. Figure 12 shows examples of scenes generated by participants. After generating the scene, we demonstrated how participants could navigate our interface. They then completed a survey about the usability of the scanning interface.
We use the participant-generated SplatOverflow scene in part 2 of our study. If the user’s input failed to generate a scene, they were given a placeholder scene featuring the same hardware.
@@ 8.3. Part 2: Following Instructions in SplatOverflow
Figure 13. This figure is a chart displaying the results of a Likert scale question. The questions ask how well users understood the problem they were fixing, how well they understood the instructions on how to fix the problem and how well they feel they did at carrying out the instructions. The chart repeats these questions for two issues the users fixed. The chat indicates most people felt positively about all questions in both conditions.
In part 2, we presented the participant with two hardware issues in random order and asked them to follow instructions rendered in SplatOverflow to remedy the issues. The issues were: correcting the belt tension in the printer’s x-axis (issue 1) and clearing a jam in the extruder (issue 2). These issues were selected as they are commonly occurring issues with this printer and have a variety of online tutorials dedicated to them (Repair, 2019; 3D, 2023). The order of the issues was randomized, and participants were asked to follow the instructions to the best of their ability. For both issues, 91.7% of participants could successfully follow instructions in SplatOverflow to remedy the problem. Following each issue, we asked participants to answer questions about their experience using SplatOverflow to troubleshoot the issue. Figure 13 shows participants’ responses on a Likert scale. Overall, users found that instructions rendered in SplatOverflow were easy to follow and well-aligned with the scanned hardware.
@@ 8.4. System Usability
After completing both parts of the study, participants evaluated the overall usability of the system through a survey. We use the System Usability Scale (Brooke, [n. d.]) and our system scored 86.25 / 100.
In addition, participants shared insights on how SplatOverflow could be expanded to assist with troubleshooting in other contexts. Participants felt that SplatOverflow would be helpful when facing issues with unfamiliar hardware: ”It is very helpful for me to use and repair, especially when I’m not familiar with the machine itself” (Participant 1). They also compared our system to existing methods of troubleshooting: ”I was working with robots (Turtlebot 4 setup). Their documentation was too wordy, and it was hard to locate the solution from that. So, we had to post GitHub issues and wait for a response. It was sometime hard to explain them my specific problem. Of course, if there was a system that I could have scanned my problem and send them, it would have saved a lot of time” (Participant 3).
@@ 9.1.1. Hardware Support for End-Users
SplatOverflow offers a new way to collaborate on hardware. For workflows such as assembly, routine maintenance, operating instructions, and troubleshooting, SplatOverflow provides a novel way for users to receive guidance from collaborators or retrieve associated documentation. By connecting physical instances of hardware to their CAD model and associated documentation, end-users can inspect the ’source’ of the hardware they are working with. This direct connection between the physical part and documentation provides a more direct way to navigate documentation that can exist in disparate locations.
Moreover, users can seek help through SplatOverflow, analogous to asking for help in a forum. In software, platforms such as GitHub (Dabbish et al., 2012) and StackOverflow (Anderson et al., 2012) have been essential for helping users learn about new software, troubleshoot issues, and scale maintenance efforts to support distributed projects. By supporting asynchronous communication about physical things, SplatOverflow presents a similar workflow for hardware.